Because printing is different from typewriting, there are a number of things that you have to do differently when preparing an input file than if you were just typing the document directly. Quotation marks like “this” have to be handled specially, as do quotes within quotes: “ ‘this’ is what I just wrote, not ‘that’ ”.

Dashes come in three sizes: an intra-word dash, a medium dash for number ranges like 1–2, and a punctuation dash—like this.

A sentence-ending space should be larger than the space between words within a sentence. You sometimes have to type special commands in conjunction with punctuation characters to get this right, as in the following sentence. Gnats, gnus, etc. all begin with G. You should check the spaces after periods when reading your output to make sure you haven’t forgotten any special cases. Generating an ellipsis … with the right spacing around the periods requires a special command.

L^{A}TEX interprets some common characters as commands, so you must type special commands to generate them. These characters include the following: $ & % # { and }.

In printing, text is usually emphasized with an italic type style.

A long segment of text can also be emphasized in this way. Text within such a segment can be given additional emphasis.

It is sometimes necessary to prevent L^{A}TEX from breaking a line where it might otherwise do so. This may be at a space, as between the “Mr.” and “Jones” in “Mr. Jones”, or within a word—especially when the word is a symbol like itemnum that makes little sense when hyphenated across lines.

Footnotes^{1} pose no problem.

L^{A}TEX is good at typesetting mathematical formulas like x − 3y + z = 7 or a_{1} > x^{2n} + y^{2n} > x′ or (A, B) = ∑_{i} a_{i}b_{i}. The spaces you type in a formula are ignored. Remember that a letter like x is a formula when it denotes a mathematical symbol, and it should be typed as one.

2 Displayed Text

Text is displayed by indenting it from the left margin. Quotations are commonly displayed. There are short quotations

This is a short a quotation. It consists of a single paragraph of text. See how it is formatted.

and longer ones.

This is a longer quotation. It consists of two paragraphs of text, neither of which are particularly interesting.

This is the second paragraph of the quotation. It is just as dull as the first paragraph.

Another frequently-displayed structure is a list. The following is an example of an itemized list.

• This is the first item of an itemized list. Each item in the list is marked with a “tick”. You don’t have to worry about what kind of tick mark is used.

3

1This is an example of a footnote.